OpenStudy (anonymous):

PLEASE HELP ME Solve -2x^2+3x-9=0 a. x=(3+/-3i)sqr7/4 b. x=(3+/-9i)sqrt7/4 c. x=(-3+/-3i)sqrt7/4 d. x=-3+/-9i)sqrt7/4

OpenStudy (anonymous):

{"x = 3/4 (1-i sqrt(7)) ~~ 0.75-1.98431 i", "x = 3/4 (1+i sqrt(7)) ~~ \ 0.75+1.98431 i"}

OpenStudy (anonymous):

Thats not right..

OpenStudy (anonymous):

Divide both sides by -2: x^2-(3 x)/2+9/2 = 0 Subtract 9/2 from both sides: x^2-(3 x)/2 = -9/2 Add 9/16 to both sides: x^2-(3 x)/2+9/16 = -63/16 Factor the left hand side: (x-3/4)^2 = -63/16 Take the square root of both sides: sqrt(x-3/4)^2 = (3 i sqrt(7))/4 Eliminate the absolute value: x-3/4 = -(3 i sqrt(7))/4 or x-3/4 = (3 i sqrt(7))/4 Add 3/4 to both sides: x = 3/4 (1-i sqrt(7)) or x-3/4 = (3 i sqrt(7))/4 Add 3/4 to both sides: x = 3/4 (1-i sqrt(7)) or x = 3/4 (1+i sqrt(7))